OpenStudy (anonymous):

find the equation of the locus of a point that moves so that its distance from the line 3x-4y+9=0 is always 5

ganeshie8 (ganeshie8):

let that point be P (x, y)

ganeshie8 (ganeshie8):

distance from a point to a line = \(\frac{Ax+By+C}{\sqrt{A^2 + B^2}}\)

ganeshie8 (ganeshie8):

can you put the equation for distance ? from point P(x, y) to line 3x-4y+9=0

OpenStudy (anonymous):

oh so it would be (3x-4y+9)/square root( 3^2 +(-4)^2)

OpenStudy (anonymous):

(3x -4y+9)/5 =5 3x -4y+9 =25 3x-4y -16 =0

OpenStudy (anonymous):

is that right ganeshi8?

OpenStudy (anonymous):

btw i thought there was an ABSOLUTE SIGN around Ax + By + C for the formula

ganeshie8 (ganeshie8):

yep! u r right.. we must get two lines

ganeshie8 (ganeshie8):

|(3x -4y+9)|/5 =5

ganeshie8 (ganeshie8):

3x-4y+9 = 25, and, 3x-4y + 9 = -25

OpenStudy (anonymous):

ngawww i get it !

ganeshie8 (ganeshie8):

it tricked me.... i was thinking there should be two lines at the very end. nice catch :)

OpenStudy (anonymous):

thanks again - LOL and about writing GAPS yesterday, it wasn't meant to be rude or anything, just expressing my gratitude for your help ;)

ganeshie8 (ganeshie8):

lol fine i get it nw :))

ganeshie8 (ganeshie8):